The closest distance a book can be read from a pair of reading eyeglasses (Power = 1.54 dp) is 27.5 cm. What is the near distance?(Assume a distance between the eyeglasses and the eyes to be 2.50 cm) please show work clearly and explain. i am used to variables Di and Do if you use other variables like s or v or p, PLEASE explain what those variables represent! thank u!
Answer #1

Given:

  • Power of eyeglasses (P) = 1.54 D

  • Distance from eyeglasses to book (Do) = 27.5 cm

  • Distance from eyeglasses to eyes (d) = 2.50 cm


Goal:

Find the near distance (Dnear), which is the closest distance the eye can focus without the glasses.


Variables:

  • Do: Object distance (book to glasses) = 27.5 cm

  • Di: Image distance (glasses to eye's near point)

  • Dnear: Near point distance (eye's natural closest focus point)

  • d: Distance from glasses to eye = 2.50 cm

Steps:

  1. Convert distances to meters (since Power is in diopters, D):

    • Do=27.5cm=0.275m

    • d=2.50cm=0.025m

  2. Use the lens formula to find the image distance (Di):

    P=1Di1Do

    Rearrange to solve for Di:

    1Di=P+1Do=1.54+10.275=1.54+3.636=5.176DDi=15.1760.1932m=19.32cm

  3. Relate the image distance to the near point (Dnear):
    The image formed by the glasses acts as the object for the eye. Since the eye is 2.50 cm from the glasses:

    Dnear=Di+d=19.32cm+2.50cm=21.82cm

Answer:

The near distance (closest distance the eye can focus without glasses) is 21.8 cm.


answered by: anonymous